Check whether V = R^2 is a vector space with respect to the operations (x1,y1) + (x2,y2) = (x1 + x2 - 2, y1 + y2 - 3) and k(x,y) = (kx + 2k -2, ky -3k + 3), k is a real number
1
Expert's answer
2020-12-06T17:24:51-0500

Answer

The above vector space V=R2 follows all vector space as follow

1.associativity properties

(a+b) +c=a+(b+c)

2.commutative law

"a+b=b+a"

There fore above V is vector space.